OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (anonymous):

do you know the relationship between gradients of two perpendicular lines?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well, they have an inverse negative relationship, i.e if gradient of line 1 is m the gradient of the line perpendicular to it is \[-1/m\]

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

don't you want to know how to do it?

OpenStudy (anonymous):

yes i'm saying.. the process

OpenStudy (anonymous):

in your case the gradient is 1/3 so (y-1) / (x--6) = 1/3 3(y-1) = x+6 3y-3 = x+6 3y = x+6-3 y = 1/3x +3